User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!

Chair/Presenting: Lucas/Arwin

Attendees: Lucas, Jonny, Tiege, Bryn, Jim, Arwin, Ellen,

Not Attending: Ross, Nick


* Elevation drives, are they fixed now? We don't know. Keep note if elevation drives don't start.

* how do you fix the monitors if they go all funky? - rebooted ops2 (keep doing it until it works)

*we're getting new monitors soon (replacing the 6 with ones from the 'lounge' area)

* seems to be a few errors with eremote control being started:

19/5 - 13:40UT - eremote control cannot be started, using pcfs vnc (Ross) - Lucas will chase this up with Ross (like what the actual problem was) 23/5 - 14:00UT - eremote control cannot be started still, using new version of eremote which appears to work (Ross) - 22:00UT - e-remotecontrol not working on ops1hb. SSH host connect OK, but no version can talk to ercd server on pcfshb (Warren)

* Ke and Yg back! They begin tomorrow (3/6/16) morning for R4742. There are VNC sessions you need to access on new PCs to check the masers in order to access the VNC sessions from ops2/4, in a terminal:

Katherine: vncviewer

Yarragadee vncviewer

The password for both is “**”.

*Hb 12m is going offline Friday for about a month, will only have 2 ant for a little bit.

* no more rebooting Field Systems.

* Have to exit the log mon at Mt Pleasant after you leave otherwise the alarms go crazy.


/home/www/auscope/opswiki/data/pages/operations/agenda2016_06_02.txt · Last modified: 2016/06/02 02:13 by Lucas Hyland